OpenStudy (anonymous):

the perimeter of a rectangle is 78m.the width is 7m less than the length. find the length AND width.

jhonyy9 (jhonyy9):

P=2l+2w and you know again that w=l-7 now hope you will can calcule it easy ,yes ?

OpenStudy (anonymous):

not exactly, but i got so far this...hang on

jhonyy9 (jhonyy9):

@AriPotta hope you can calcule it for medal ,ok ?

OpenStudy (anonymous):

78=2(l-7)+2l 782l-14+2l 78+14=92 92=4l

OpenStudy (aripotta):

that's right, just divide both sides by four to get the answer for l

OpenStudy (anonymous):

78=2(l-7)+2l 78=2l-14+2l 78+14=92 92=4l

OpenStudy (anonymous):

sorry had a mistype

OpenStudy (anonymous):

which side of what?

OpenStudy (aripotta):

92 = 4l

OpenStudy (aripotta):

when you divide both sides by four, what do you get?

OpenStudy (anonymous):

88=L

OpenStudy (anonymous):

wait no i subtratcred

OpenStudy (aripotta):

yea, you gotta divide lol

OpenStudy (anonymous):

subtracted* typo

OpenStudy (anonymous):

23=L

OpenStudy (aripotta):

yup. and since w = l - 7, and l = 23, we know that w = 23 - 7

OpenStudy (aripotta):

which would make w = ?

OpenStudy (anonymous):

for the length it is 23, but about the width, being 16?

OpenStudy (aripotta):

yea, you're right. the length is 23, and the width is 16 :)

OpenStudy (anonymous):

ahh okay!!!thanks!!!

OpenStudy (anonymous):

may i have a medal please?:)

OpenStudy (anonymous):

thanks lol:)